Fermi National Laboratory
 

Operation Summary for November 11, 2009 - November 13, 2009

 

Wednesday Nov. 11

The day shift began with Operations monitoring the Tevatron (TeV) store 7341 (114.10E30), Recycler stash (111.5E10) and the Antiproton Source stack (9.2E10), with the Linac, Booster and MiniBooNE operating normally, with PBar and NuMI off due to LAM42 repairs.

During the day shift, power supply personnel worked on LAM42 and had it back on line at 12:08 PM. Beam returned to NuMI and PBar at 12:10 PM.

During the evening shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Store 7341 collided until the scheduled termination time of 8:35 PM. Operations put the TeV into HEP shot setup at 9:07 PM and established store 7342 at 10:30 PM with an average initial luminosity (AIL) of 225.99E30.

Thursday Nov. 12

The midnight shift began with Operations monitoring store 7342 (145.5E30) and the stash (71.2E10), with Pbar stacking (23.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. Store 7242 collided until 2:23 AM, when a PBar kicker (AAK2) prefired and caused a multi house quench. The kicker was put into condition mode and the rest of the shift was spent recovering from the quench.

During the day shift, Operations delivered beam to the Pbar Source, NuMI, and MiniBooNE when available. At 9:23 AM, the Low Beta quads (DA & DB) quenched during checkout. The Tevatron completed a dry squeeze at 12:19 PM and was put into shot setup at 3:03 PM.

On the evening shift, Operations established store 7346 at 4:56 PM with an average initial luminosity of 273.20E30. At 11:52 PM, LAM61 tripped off, an expert was called to investigate, no beam to NuMI for the rest of the shift.

Friday Nov. 13

The midnight shift began with Operations monitoring store 7346 (111.02E30) and the stash (207.9E10), with Pbar stacking (28.0E10), and with MiniBooNE taking beam.

NuMI was off due to problems with LAM61.

During the midnight shift, Operations delivered beam to the Pbar Source and MiniBooNE when available. The expert repaired LAM61 and beam returned to NuMI at 1:02 AM.

The Plans for Friday and the Weekend

The plan is to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/13/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory